write the slope-intercept form of the equation of the line passing through the points (-6,-2) and parallel to this line y=-3x-4
Answer by stanbon(75887)

write the slope-intercept form of the equation of the line passing through the points (-6,-2) and parallel to this line y=-3x-4
-----------------------------
The given line has a slope of -3.
----
Any line parallel to it must have a slope of -3.
----
Equation Form:
y = mx + b
---
You know x = -6 when y = -2 and you know m = -3 ; solve for "b":
----
-2 + -3(-6)+b
b = -2-18 = -20
----
Equation:
y = -3x-20